OpenStudy (anonymous):

Given the equation 3x - 3 = 6x + 1, select the reasoning that correctly solves for x. Add 3, subtract 6x, then divide by 3. Is my choice, out of: Add 3, subtract 3x, then divide by -3. Add 3, subtract 6x, then divide by 3. Subtract 1, subtract 3x, then divide by 3. Subtract 1, subtract 6x, then divide by -3. My reasoning is: 3x - 3 = 6x + 1 3x = 6x +4 -6x -6x -3x =4? I might be wrong, I'm iffy.

OpenStudy (anonymous):

Subtract 1, subtract 3x, then divide by 3. I think this is it, makes more sense to me.

pooja195 (pooja195):

The answer you got is incorrect

OpenStudy (anonymous):

Both of them?

pooja195 (pooja195):

Yep

pooja195 (pooja195):

Hint: We only have 3 steps Add Subtract Divide ^-----these are not listed in order

OpenStudy (anonymous):

Add 3, subtract 3x, then divide by -3. x-x

pooja195 (pooja195):

;-; no Thats not even one of your options

pooja195 (pooja195):

How did u get subtract 3x?

OpenStudy (anonymous):

o-o It's the first option.

pooja195 (pooja195):

Ohhhhh oops u listed 2 of the same options

OpenStudy (anonymous):

No, I just showed the answer I chose then after I listed all of the choices where they came from. Subtract 1, subtract 3x, then divide by 3. 3x - 3 = 6x + 1 3x -4 = 6x -4 = 3x Not it ----------- Add 3, subtract 6x, then divide by 3. 3x - 3 = 6x + 1 3x = 6x +4 -3x = 4 Not it. -------- Subtract 1, subtract 6x, then divide by -3. 3x - 3 = 6x + 1 3x - 4 = 6x -3x -4 No. ------- Add 3, subtract 3x, then divide by -3. 3x - 3 = 6x + 1 3x = 6x +4 6x + 4 No? ---- I'm confused. I solve these problems a bit differently so this threw me off.

OpenStudy (anonymous):

Like, 3x - 3 = 6x + 1. I usually subtract the variable. 3x - 3 = 6x + 1 -3x -3x Leaves me with -3 = 3x +1 -4 = 3x So... Idk.

pooja195 (pooja195):

Subtract 6x from both sides. 3x−3−6x=6x+1−6x −3x−3=1 Add 3 to both sides. −3x−3+3=1+3 −3x=4 Divide both sides by -3. x=-4/3

OpenStudy (anonymous):

That isn't one of the choices.

pooja195 (pooja195):

−3x=4 is one of the steps i assume thats what they want in the answer

OpenStudy (anonymous):

It would be if the "divide by -3" part would be right if it was a +3. This is why I'm confused.

OpenStudy (anonymous):

Wouldn't I be able to subtract 1 instead as well? That would fit the final choice I think.

pooja195 (pooja195):

I'm confused prehaps @KendrickLamar2014 or @jabez177 can help?

OpenStudy (kendricklamar2014):

So, what may I be of assistance?

OpenStudy (anonymous):

The question I posted. None of the options give any sense at all, and when we present one that is true, it is not an answer choice. My closest guess is: Subtract 1, subtract 3x, then divide by 3.

OpenStudy (kendricklamar2014):

None of the answers make sense to me :(

OpenStudy (anonymous):

One gives the coefficient on the wrong side, and the other two give the number and variable on the same side, which makes it impossible to isolate the variable.

jabez177 (jabez177):

I'm sorry but Math isn't my best subject. If you need Assistance in English, Science, or History, feel free to tag me. :)

pooja195 (pooja195):

@IrishBoy123 ?

OpenStudy (anonymous):

That's what's confusing us. I worked around 15 minutes on that question and I can't seem to make sense of any of it. I'm just about to choose a random answer and link my teacher to this thread so she knows you guys can't solve it either, and you're all 99's and qualified helpers. If anything, it has to be a typo. There's no other way to explain it.

OpenStudy (anonymous):

1, subtract 3x, then divide by 3. Is the correct answer. I did the assignment and got a 100%.